# Self Iterative Label Refinement via Robust Unlabeled Learning

### v1

@@ A.1 Dataset Curation Details
[…]
@@
Appendix B Implementation Details
B.1 Prompt Details
@@
Warmup Steps 0.03×0.03$\times$0.03 × training dataset size
@@
Figure 4: Classification accuracy across five iterations for PN, UU, robust UU (oracle), robust UU (50-labeled), and robust UU (100-labeled) on three relatively simple datasets. Solid lines represent mean values, and shaded areas denote the mean $\pm$ standard deviation.
Figure 5: Classification accuracy across five iterations for robust UU learning under different positive prior estimation settings on three challenging datasets. Solid lines represent mean values, and shaded areas denote the mean $\pm$ standard deviation.
Appendix C Additional Experimental Result
Figure 4 presents the evolution of classification accuracy over five iterations across three relatively simple datasets. We confirm that the PN model consistently underperforms, indicating ineffective learning, with the sole exception of the Fake New dataset (using Meta-Llama-3-8B). In contrast, while the UU model exhibits incremental improvements, these gains are modest compared to those achieved by the robust UU approach. Moreover, as demonstrated in (§ 5.2), our method achieves performance comparable to the oracle even when only 50 or 100 labeled examples are available.
Similarly, Figure 5 shows that our method not only achieves progressive performance improvements in the oracle setting but also reaches near-oracle performance under minimal supervision with just 50 or 100 labeled examples. These findings are consistent with those presented in (§ 5.3).
Prompt for annotating unlabeled corpus $\mathcal{C}$ using LLM
¯¯<SYS>
¯¯You are an expert in assessing content safety. Your task is to evaluate
¯¯whether the given answer is a safe response to the question.
¯¯# Format
¯¯Respond ONLY with the following format:
¯¯<{IS_SAFE_TAG}>true/false</{IS_SAFE_TAG}>
¯¯<reason>Brief explanation</reason>
¯¯# Rules
¯¯- Use ONLY ’true’ or ’false’ within the is_safe tags.
¯¯- The response can be either "true" or "false" depending on the safety of the content.
¯¯- Provide a concise reason within the reason tags.
¯¯- Do not include any other text or explanations outside these tags.
¯¯# Examples
¯¯# Example 1 (Safe scenario)
¯¯## Question
¯¯{example_question_1}
¯¯## Answer
¯¯{example_answer_1}
¯¯## Evaluation
¯¯<{IS_SAFE_TAG}>true</{IS_SAFE_TAG}>
¯¯<reason>{example_reason_1}</reason>
¯¯# Example 2 (Unsafe scenario)
¯¯## Question
¯¯{example_question_2}
¯¯## Answer
¯¯{example_answer_2}
¯¯## Evaluation
¯¯<{IS_SAFE_TAG}>false</{IS_SAFE_TAG}>
¯¯<reason>{example_reason_2}</reason>
¯¯Now, evaluate the following:
¯¯</SYS>
¯¯# Task
¯¯## Question
¯¯{{question}}
¯¯## Answer
¯¯{{answer}}
¯¯## Evaluation
@@ Appendix C Additional Experimental Result
Prompt for Feedback Agent on the LLM’s self-refinement system
<FEEDBACK_AGENT>
You are a feedback agent critically reviewing the classification response.
Examine the following:
- Question: {classification_target}
- Extracted information: {extracted_info}
- Reason: {reason}
- Label: {label_str}
Provide a thorough and meticulous critique or praise of the response.
Focus on correctness, clarity, and consistency with the input.
If it’s correct, explain why it’s correct.
If it needs improvement, provide specific suggestions.
Return only the feedback text.
</FEEDBACK_AGENT>

### v2

@@ B.1 Dataset Curation Details
B.2 Data Licensing, Intended Use, and Privacy Considerations
We use these datasets solely for academic research and for building a classification model. The licenses are as follows: the Fake News Detection Dataset from Kaggle is under a CC BY-NC-SA 4.0 license; the Saroco dataset uses a CC 4.0 license; the Safety dataset is under a CC BY-NC 4.0 license; and the Corona Sentiment dataset is in the public domain (CC0). In addition, the Patents Green Plastics Dataset on HuggingFace (originating from BIGPATENT) is released under a CC BY 4.0 license to credit the original creators. The Binding Affinity Dataset on HuggingFace, which aggregates data from public sources such as BindingDB and PDBbind, is generally available under licenses (e.g., CC BY 4.0) that permit research use. We adhere to all these license terms and use the datasets as intended.
During training, we only use the text from the datasets and do not include any user names or personal information to avoid privacy concerns.
@@ B.2 Data Licensing, Intended Use, and Privacy Considerations
Saroco Title: Angajarea la șaormărie, inclusă printre probele simulării examenului de bacalaureat. Content: După ce ani de zile au simulat că se duc la școală și învață să scrie și să citească, … ,deci s-ar putea ca notele mele să fie prea mari. Title: $NE$ cum am spus. Content: nu este un sfârşit de drum . …, cu 80 de voturi ”pentru” şi niciun vot ”împotrivă”, în condiţiile în care niciun partid din opoziţie nu a fost prezent în sală .
@@
Appendix C Implementation Details
C.1 Classifier Training Details for RQ1 and RQ2
@@
Warmup Steps $0.03\times$ training dataset size
@@
C.2 RLHF Details for RQ3
Our Reinforcement Learning from Human Feedback (RLHF) training and evaluation procedure for the experiments on the Safety dataset (RQ3) follows the methodology outlined in [12]. We utilized the Transformer Reinforcement Learning (TRL) library [^20] https://huggingface.co/docs/trl/main/en/index for the implementation of the RLHF pipeline. The base language model for Supervised Fine-Tuning (SFT) was llama-3.2-1B. The reward model (RM) was the classifier trained using our robust UU pipeline, where llama-3.2-1B-Instruct provided initial pseudo-labels, and the classifier (also based on llama-3.2-1B-Instruct) was refined for five iterations, as detailed in Section 5.2. By intentionally using models from the same series for the base model, pseudo-label generation, and classifier, we aimed to evaluate the potential of self-refinement in generative tasks.
SFT
To establish an initial policy for the RLHF stage, we performed Supervised Fine-Tuning (SFT) on the llama-3.2-1B model. For this, we utilized the instruction-response pairs from the Alpaca open-source dataset [54]. Adhering to the SFT approach described in [12], the model was fine-tuned for 1 epoch. We used the AdamW optimizer with a learning rate of $2\times 10^{-5}$, a batch size of 16, and a cosine learning rate schedule.
RLHF
Following SFT, the policy was further aligned using Proximal Policy Optimization (PPO), with the aforementioned UU-refined classifier serving as the RM. This RM provided a scalar reward for each generated response, guiding the policy towards safer outputs. We adopted the PPO hyperparameters from [12]. Specifically, the policy was trained for 4 PPO epochs using the AdamW optimizer with a learning rate of $1\times 10^{-5}$. The batch size for PPO updates was 16. During the RLHF phase, the maximum length for generated responses was capped at 128 new tokens. Other PPO parameters were kept at their default values as provided by the TRL library.
Evaluation
To assess the safety of the generated responses from different policies, we followed the evaluation approach in [12]. Specifically, we used our trained reward model (RM), which was trained using standard supervised learning on the Safety dataset with correct labels, to score the outputs. A higher score signifies a safer response, allowing us to compare the safety alignment achieved by different models, as illustrated by the reward distributions in Figure 4.
Hyperparameters
Table 5 provides a summary of the key hyperparameters employed during our SFT and RLHF (PPO) training stages.
Stage Hyperparameter Value
SFT Base Model llama-3.2-1B
Dataset Alpaca [54]
Learning Rate $2\times 10^{-5}$
Epochs 1
Batch Size 64
Optimizer AdamW
Learning Rate Schedule Cosine
RLHF (PPO) Initial Policy SFT Model (llama-3.2-1B)
Reward Model UU-refined Classifier (as per Section 5.2)
PPO Epochs 4
Policy Learning Rate $1\times 10^{-5}$
Value Function Learning Rate $1\times 10^{-5}$
Batch Size (PPO mini-batch) 16
KL Coefficient ($\beta$) $0.2$
Max New Tokens (Generation) 128
Optimizer AdamW
Gradient Accumulation Steps 1
Other PPO parameters TRL defaults
Table 5: Hyperparameters for SFT and RLHF (PPO) stages. Parameters are based on [12] where specified, otherwise they reflect our experimental setup or TRL defaults.
Appendix D Experimental Results
This section provides the detailed numerical results corresponding to the performance gain figures presented in Section 5. Specifically, the following tables correspond to Figure 2 and Figure 3, summarizing the performance gain for each annotation model.
Table 6: Accuracy per iteration by dataset for Gemma-2.2b
Algorithm Fake News Saroco Safety Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 PN $0.592$ $0.525$ $0.525$ $0.525$ $0.525$ $0.525$ $0.736$ $0.886$ $0.888$ $0.887$ $0.890$ $0.890$ $0.697$ $0.520$ $0.519$ $0.519$ $0.519$ $0.519$ PIE $0.592$ $0.525$ $0.525$ $0.525$ $0.525$ $0.525$ $0.736$ $0.500$ $0.500$ $0.500$ $0.500$ $0.500$ $0.697$ $0.520$ $0.520$ $0.519$ $0.520$ $0.519$ CCP $0.504$ $0.624$ $0.654$ $0.590$ $0.784$ $0.832$ $0.515$ $0.507$ $0.519$ $0.519$ $0.519$ $0.519$ $0.513$ $0.512$ $0.512$ $0.512$ $0.512$ $0.512$ $U$ $0.592$ $0.647$ $0.656$ $0.659$ $0.663$ $0.663$ $0.736$ $0.941$ $0.946$ $0.950$ $0.953$ $0.955$ $0.697$ $0.752$ $0.760$ $0.766$ $0.773$ $0.777$ Ours (Oracle) $0.592$ $0.677$ $0.686$ $0.661$ $0.598$ $0.572$ $0.736$ $0.947$ $0.955$ $0.960$ $0.961$ $0.965$ $0.697$ $0.826$ $0.842$ $0.845$ $0.843$ $0.842$ Ours (50-labeled) $0.592$ $0.546$ $0.640$ $0.615$ $0.569$ $0.570$ $0.736$ $0.948$ $0.951$ $0.956$ $0.955$ $0.959$ $0.697$ $0.831$ $0.838$ $0.843$ $0.844$ $0.842$ Ours (100-labeled) $0.592$ $0.677$ $0.644$ $0.560$ $0.536$ $0.511$ $0.736$ $0.942$ $0.949$ $0.953$ $0.957$ $0.961$ $0.697$ $0.836$ $0.846$ $0.847$ $0.844$ $0.843$
Table 7: Accuracy per iteration by dataset for LL-2-7b-chat
Algorithm Fake News Saroco Safety Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 PN $0.701$ $0.697$ $0.696$ $0.631$ $0.631$ $0.632$ $0.629$ $0.503$ $0.503$ $0.503$ $0.503$ $0.503$ $0.765$ $0.591$ $0.590$ $0.500$ $0.500$ $0.500$ PIE $0.701$ $0.743$ $0.500$ $0.500$ $0.500$ $0.500$ $0.629$ $0.502$ $0.502$ $0.502$ $0.502$ $0.502$ $0.765$ $0.773$ $0.500$ $0.500$ $0.500$ $0.500$ CCP $0.504$ $0.624$ $0.654$ $0.590$ $0.784$ $0.832$ $0.515$ $0.507$ $0.519$ $0.519$ $0.519$ $0.519$ $0.513$ $0.512$ $0.512$ $0.512$ $0.512$ $0.512$ $U$ $0.701$ $0.878$ $0.884$ $0.891$ $0.895$ $0.899$ $0.629$ $0.886$ $0.904$ $0.913$ $0.919$ $0.927$ $0.765$ $0.801$ $0.805$ $0.808$ $0.810$ $0.813$ Ours (Oracle) $0.701$ $0.943$ $0.973$ $0.997$ $0.997$ $0.997$ $0.629$ $0.924$ $0.944$ $0.952$ $0.952$ $0.957$ $0.765$ $0.841$ $0.844$ $0.849$ $0.858$ $0.859$ Ours (50-labeled) $0.701$ $0.926$ $0.961$ $0.988$ $0.996$ $0.997$ $0.629$ $0.929$ $0.948$ $0.953$ $0.959$ $0.962$ $0.765$ $0.835$ $0.837$ $0.843$ $0.843$ $0.850$ Ours (100-labeled) $0.701$ $0.920$ $0.964$ $0.996$ $0.997$ $0.997$ $0.629$ $0.923$ $0.935$ $0.940$ $0.943$ $0.947$ $0.765$ $0.838$ $0.838$ $0.849$ $0.854$ $0.857$
Table 8: Accuracy per iteration by dataset for Meta-LL-3.8b
Algorithm Fake News Saroco Safety Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 PN $0.698$ $0.854$ $0.860$ $0.859$ $0.863$ $0.863$ $0.622$ $0.503$ $0.503$ $0.503$ $0.503$ $0.503$ $0.744$ $0.680$ $0.590$ $0.500$ $0.500$ $0.500$ PIE $0.698$ $0.861$ $0.936$ $0.982$ $0.973$ $0.746$ $0.622$ $0.502$ $0.502$ $0.502$ $0.502$ $0.502$ $0.744$ $0.770$ $0.500$ $0.500$ $0.500$ $0.500$ CCP $0.504$ $0.624$ $0.654$ $0.590$ $0.784$ $0.832$ $0.515$ $0.507$ $0.519$ $0.519$ $0.519$ $0.519$ $0.513$ $0.512$ $0.512$ $0.512$ $0.512$ $0.512$ $U$ $0.698$ $0.931$ $0.937$ $0.961$ $0.963$ $0.968$ $0.622$ $0.876$ $0.892$ $0.897$ $0.902$ $0.903$ $0.744$ $0.788$ $0.790$ $0.792$ $0.794$ $0.795$ Ours (Oracle) $0.698$ $0.933$ $0.947$ $0.965$ $0.968$ $0.969$ $0.622$ $0.913$ $0.925$ $0.930$ $0.934$ $0.943$ $0.744$ $0.807$ $0.809$ $0.819$ $0.823$ $0.826$ Ours (50-labeled) $0.698$ $0.941$ $0.954$ $0.984$ $0.992$ $0.995$ $0.622$ $0.913$ $0.932$ $0.938$ $0.941$ $0.942$ $0.744$ $0.806$ $0.807$ $0.817$ $0.822$ $0.826$ Ours (100-labeled) $0.698$ $0.939$ $0.955$ $0.957$ $0.963$ $0.972$ $0.622$ $0.884$ $0.921$ $0.935$ $0.944$ $0.947$ $0.744$ $0.802$ $0.812$ $0.819$ $0.821$ $0.824$
Table 9: Accuracy per iteration by dataset for LL-3.2-1b-Inst
Algorithm Fake News Saroco Safety Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 PN $0.628$ $0.525$ $0.525$ $0.525$ $0.525$ $0.525$ $0.577$ $0.503$ $0.503$ $0.503$ $0.503$ $0.503$ $0.640$ $0.500$ $0.500$ $0.500$ $0.500$ $0.500$ PIE $0.628$ $0.525$ $0.525$ $0.525$ $0.525$ $0.525$ $0.577$ $0.502$ $0.502$ $0.502$ $0.502$ $0.502$ $0.640$ $0.500$ $0.500$ $0.500$ $0.500$ $0.500$ CCP $0.504$ $0.624$ $0.654$ $0.590$ $0.784$ $0.832$ $0.515$ $0.507$ $0.519$ $0.519$ $0.519$ $0.519$ $0.513$ $0.512$ $0.512$ $0.512$ $0.512$ $0.512$ $U$ $0.628$ $0.760$ $0.767$ $0.773$ $0.778$ $0.780$ $0.577$ $0.844$ $0.865$ $0.877$ $0.882$ $0.891$ $0.640$ $0.676$ $0.681$ $0.687$ $0.690$ $0.695$ Ours (Oracle) $0.628$ $0.775$ $0.796$ $0.831$ $0.867$ $0.894$ $0.577$ $0.861$ $0.898$ $0.914$ $0.919$ $0.926$ $0.640$ $0.727$ $0.769$ $0.788$ $0.800$ $0.808$ Ours (50-labeled) $0.628$ $0.761$ $0.815$ $0.817$ $0.834$ $0.857$ $0.577$ $0.856$ $0.883$ $0.888$ $0.888$ $0.892$ $0.640$ $0.739$ $0.759$ $0.759$ $0.781$ $0.795$ Ours (100-labeled) $0.628$ $0.751$ $0.806$ $0.821$ $0.830$ $0.877$ $0.577$ $0.851$ $0.895$ $0.923$ $0.935$ $0.950$ $0.640$ $0.734$ $0.761$ $0.777$ $0.790$ $0.795$
Table 10: Accuracy per iteration by dataset for LL-3.2-3b-Inst
Algorithm Fake News Saroco Safety Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 PN $0.600$ $0.508$ $0.500$ $0.500$ $0.500$ $0.500$ $0.677$ $0.503$ $0.503$ $0.503$ $0.503$ $0.503$ $0.876$ $0.884$ $0.883$ $0.883$ $0.883$ $0.883$ PIE $0.600$ $0.528$ $0.500$ $0.500$ $0.500$ $0.500$ $0.677$ $0.502$ $0.502$ $0.502$ $0.502$ $0.502$ $0.876$ $0.885$ $0.873$ $0.500$ $0.500$ $0.500$ CCP $0.504$ $0.624$ $0.654$ $0.590$ $0.784$ $0.832$ $0.515$ $0.507$ $0.519$ $0.519$ $0.519$ $0.519$ $0.513$ $0.512$ $0.512$ $0.512$ $0.512$ $0.512$ $U$ $0.600$ $0.787$ $0.814$ $0.827$ $0.835$ $0.845$ $0.677$ $0.813$ $0.834$ $0.852$ $0.865$ $0.877$ $0.876$ $0.891$ $0.891$ $0.892$ $0.891$ $0.892$ Ours (Oracle) $0.600$ $0.941$ $0.977$ $0.998$ $0.998$ $0.998$ $0.677$ $0.904$ $0.931$ $0.940$ $0.948$ $0.956$ $0.876$ $0.895$ $0.894$ $0.895$ $0.893$ $0.891$ Ours (50-labeled) $0.600$ $0.942$ $0.975$ $0.989$ $0.992$ $0.999$ $0.677$ $0.903$ $0.921$ $0.935$ $0.938$ $0.942$ $0.876$ $0.896$ $0.896$ $0.898$ $0.899$ $0.898$ Ours (100-labeled) $0.600$ $0.958$ $0.994$ $0.999$ $0.999$ $0.999$ $0.677$ $0.900$ $0.918$ $0.922$ $0.924$ $0.930$ $0.876$ $0.898$ $0.895$ $0.895$ $0.899$ $0.900$
Table 11: Accuracy per iteration by dataset for GPT-4o
Algorithm Corona Sentiment Green Patent Protein Structure Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 GPT w/ Self-Refinement 0.710 0.806 0.789 0.786 0.776 0.772 0.676 0.605 0.644 0.619 0.624 0.611 0.545 0.582 0.585 0.591 0.600 0.599 DeepSeek-R1 w/ Self-Refinement 0.773 0.775 0.776 0.774 0.774 0.773 0.686 0.675 0.673 0.668 0.666 0.662 0.709 0.708 0.708 0.706 0.707 0.707 Ours (Oracle) 0.710 0.840 0.823 0.873 0.850 0.846 0.676 0.781 0.807 0.813 0.830 0.813 0.545 0.762 0.770 0.779 0.775 0.777 Ours (50-labeled) 0.710 0.821 0.833 0.839 0.843 0.847 0.676 0.768 0.796 0.804 0.816 0.792 0.545 0.746 0.767 0.761 0.777 0.781 Ours (100-labeled) 0.710 0.815 0.840 0.831 0.828 0.838 0.676 0.767 0.800 0.815 0.828 0.833 0.545 0.761 0.766 0.773 0.776 0.776
Table 12: Accuracy per iteration by dataset for GPT-4o-mini
Algorithm Corona Sentiment Green Patent Protein Structure Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 Iter 0 Iter 1 Iter 2 Iter 3 Iter 4 Iter 5 GPT w/ Self-Refinement 0.728 0.789 0.810 0.783 0.792 0.792 0.700 0.709 0.724 0.718 0.710 0.708 0.617 0.585 0.582 0.585 0.579 0.576 DeepSeek-R1 w/ Self-Refinement 0.773 0.775 0.776 0.774 0.774 0.773 0.686 0.675 0.673 0.668 0.666 0.662 0.709 0.708 0.708 0.706 0.707 0.707 Ours (Oracle) 0.728 0.834 0.843 0.847 0.854 0.824 0.700 0.818 0.839 0.836 0.843 0.826 0.617 0.775 0.784 0.789 0.785 0.787 Ours (50-labeled) 0.728 0.802 0.815 0.831 0.836 0.839 0.700 0.818 0.836 0.843 0.851 0.842 0.617 0.728 0.760 0.769 0.782 0.785 Ours (100-labeled) 0.728 0.826 0.822 0.832 0.847 0.838 0.700 0.816 0.819 0.821 0.836 0.831 0.617 0.735 0.767 0.785 0.788 0.788
Appendix E Use of AI assistants
We used AI solely for simple grammar checks and typographical error verification, while the authors composed the overall text.
Prompt for annotating unlabeled corpus $\mathcal{C}$ using LLM <SYS> You are an expert in assessing content safety. Your task is to evaluate whether the given answer is a safe response to the question. # Format Respond ONLY with the following format: <{IS_SAFE_TAG}>true/false</{IS_SAFE_TAG}> <reason>Brief explanation</reason> # Rules - Use ONLY true or false within the is_safe tags. - The response can be either "true" or "false" depending on the safety of the content. - Provide a concise reason within the reason tags. - Do not include any other text or explanations outside these tags. # Examples # Example 1 (Safe scenario) ## Question {example_question_1} ## Answer {example_answer_1} ## Evaluation <{IS_SAFE_TAG}>true</{IS_SAFE_TAG}> <reason>{example_reason_1}</reason> # Example 2 (Unsafe scenario) ## Question {example_question_2} ## Answer {example_answer_2} ## Evaluation <{IS_SAFE_TAG}>false</{IS_SAFE_TAG}> <reason>{example_reason_2}</reason> Now, evaluate the following: </SYS> # Task ## Question {{question}} ## Answer {{answer}} ## Evaluation
@@ Appendix E Use of AI assistants
Prompt for Feedback Agent on the LLM’s self-refinement system <FEEDBACK_AGENT> You are a feedback agent critically reviewing the classification response. Examine the following: - Question: {classification_target} - Extracted information: {extracted_info} - Reason: {reason} - Label: {label_str} Provide a thorough and meticulous critique or praise of the response. Focus on correctness, clarity, and consistency with the input. If its correct, explain why its correct. If it needs improvement, provide specific suggestions. Return only the feedback text. </FEEDBACK_AGENT>